Consider the titration of 40.00 mL of 0.500 M NaIO (sodium hypoiodite) with 0.750 M HNO3. What volume of HNO3 will give pH = pKa?
1
Expert's answer
2019-03-08T05:00:38-0500

Solution.

Firstly, we assume that the ionization constant of nitric acid is equal to one. Therefore, the number is equal to one. And this means that the acid will be in excess.

pH = -lg[H+]

lg[H+] = -1

[H+] = 0.1

n(NaIO) = 0.04 liters * 0.500 M = 0.02 moles

x = liters of HNO3

n((H+) in excess) = x*0.750-0.02

V(total) = 0.04 + x

[H+] = (n((H+) in excess))/V(total)

0.1 = (x*0.750-0.02)/(0.04+x)

x = 0.036 liters

x = 36.00 mL

Answer:

36.00 mL